OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 15 and b = 5.

OpenStudy (anonymous):

\[P=2(a+b)\] \[15=2(a+5)\] \[\frac{ 15 }{ 2 }=\frac{ 2(a+5) }{ 2 }\] \[7.5=a+5\] \[\frac{ 7.5 }{ -5 }=\frac{ a+5 }{ -5 }\]

OpenStudy (anonymous):

\[a=2.5\]